OpenStudy (anonymous):

In solving the equation (x + 4)(x – 4) = 33, Eric stated that the solution would be x + 4 = 33 => x = 29 or (x – 4) = 33 => x = 37 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning.

OpenStudy (anonymous):

you cant take them separately like that as it is a quadratic equation

OpenStudy (anonymous):

(x + 4)(x – 4) = 33 x^2-16=33 x^2=33+16 x^2=49 x=7

OpenStudy (anonymous):

thank you